Circular No. 8888

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V2467 CYGNI R. J. Russell, R. J. Rudy, D. K. Lynch, The Aerospace Corporation;

and C. E. Woodward, University of Minnesota, report 0.8- to 5.5-μm spectroscopy of V2467 Cyg on Oct. 10 and 11 UT using the Infrared Telescope Facility (+ SpeX). The spectrum of the nova has increased in excitation since coronal lines were reported last May (IAUC 8848). Lines of [Si X] and [S XI] are now seen, indicating the presence of ionizing photons with energies up to at least 447 electron volts. The rarely seen [P VII] feature at 1.37 μm is also observed. The hydrogen/helium emission-line profiles remained peaked, whereas many of the coronal lines had flatter, castellated tops. Line widths (FWHM) were ∼ 2000 km/s. The lines of neutral carbon and nitrogen were gone, but the O I lines are still present and suggest a reddening of E(B −V) = 1.7, similar to the value of 1.5 reported by Mazuk et al. (IAUC 8848). There is no indication of dust formation in the nova ejecta.

COMETS C/2007 L7–L11 (SOHO) Further to IAUC 8883, additional Kreutz sungrazing comets have been

found on SOHO website images (“discovery” observations tabulated below). C/2007 L7 was small, slightly diffuse, and very faint (mag ∼ 8). C/2007 L8 and C/2007 L9 were very diffuse (mag ∼ 7.5), the former having no tail, and the latter being elongated. C/2007 L10 was very faint and stellar in appearance in C3 images, and slightly diffuse (mag ∼ 7.5) with a hint of tail in C2 images. C/2007 L11 was stellar in appearance (mag ∼ 6.5) in C3 images, and very diffuse with a short tail in C2 images.

Comet 2007 UT α_{2000} δ_{2000} Inst. F MPEC
C/2007 L7 June 10.663 5^{h}09.^{m}7 +21^{o}06′ C2  BZ 2007-T117
C/2007 L8  10.688 5 08.7 +21 03 C2  BZ 2007-T117
C/2007 L9  10.996 5 10.0 +21 02 C2  BZ 2007-T117
C/2007 L10  13.321 5 20.9 +20 12 C3/2 BZ 2007-T122
C/2007 L11  14.821 5 27.1 +20 18 C3/2 TH 2007-T122

COMETS 192P/SHOEMAKER-LEVY AND 193P/LINEAR-NEAT Comet P/2007 T3 = P/1990 V1 (cf. IAUC 8879) has been assigned the

permanent number 192P, and comet P/2007 U2 = P/2001 Q5 (cf. IAUC 8885) has been assigned the permanent number 193P (cf. MPC 60847).

2007 November 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green